# Elasticity without a reference state: continuum mechanics of active tension nets

### v2

@@ Appendix F Kinematics of Beltrami flow
Beltrami flow induces $\mu_{\zeta|\zeta^{\prime}}=\bar{\partial}_{\zeta}{\dot{\zeta}}dt/(1+\partial_{\zeta}{\dot{\zeta}}dt)\approx\bar{\partial}_{\zeta}{\dot{\zeta}}dt=\partial_{t}\mu_{a}dt$. To find the Beltrami coefficient of the inverse of a map, one applies the chain rule to obtain the relation
@@
Now composing with $\zeta\mapsto z$ gives
@@
where the phase factor is $e^{-i2\phi_{\zeta^{\prime}|\zeta}}=(1-\overline{\partial_{\zeta}{\dot{\zeta}}})/(1-\partial_{\zeta}{\dot{\zeta}})\approx 1+i2\mathrm{Im}[\partial_{\zeta}{\dot{\zeta}}]dt$. The displacement change from $\zeta$ to $\zeta^{\prime}$ leads to an additional advective term $\mathrm{Re}[\bar{\nu}\partial_{\zeta}]\mu_{z}dt$. Taken together, one arrives at
@@
$=\partial_{t}\mu_{\zeta|z}-\partial_{t}\mu_{a}+\mathrm{Re}[\bar{v}_{(\zeta)}\partial_{\zeta}]\mu_{\zeta|z}$
@@
The final, nonlinear term can be neglected for weak anisotropy $|\mu_{\zeta|z}|\ll 1$ and captures strain-induced rotation. Figure 5 provides a geometric explanation of Eq. (76). The term $\partial_{t}\mu_{\zeta|z}$ arises from the intrinsic dynamics of $z$, in addition to the coordinate transformation $\zeta\mapsto\zeta^{\prime}$.
F.2 Variation of $|\mu_{\zeta|w}|^{2}$
@@
makes clear that the Beltrami coefficient of the new map $\zeta^{\prime}\mapsto w$ is calculated by composing $\zeta^{\prime}\mapsto\zeta\mapsto w$ [red path in Eq. (78)]. We first use that $\zeta\mapsto\zeta^{\prime}$ is simply the Beltrami flow induced by $\delta\mathbf{a}$ so $\mu_{\zeta|\zeta^{\prime}}=\bar{\partial}_{\zeta}\zeta^{\prime}/\partial_{\zeta}\zeta^{\prime}=\delta\mu_{a}$. The inverse is obtained using Eq. (74), giving
@@
Appendix G Examples
Here, we provide three concrete examples to illustrate the continuum equations derived in the main text. First, we consider two static examples. Since the static problem is mathematically equivalent to linear elasticity, these examples have no novel content, but serve to clarify the logic of the mapping from the tension manifold to the physical state. Second, we illustrate the dynamics equations by considering the example of positive tension feedback, building on the mean-field analysis in Ref. [30].
G.1 Statics: flat tension metric
As a first example, we take a uniform, incompressible, rectangular tissue patch. The set-up is sketched in Fig. 6. Let the adjacency metric be the Euclidean metric, corresponding to uniformly distributed isotropic cells. The $a$-isothermal $\zeta$ coordinates are simply Cartesian coordinates in the unit square $0\leq\zeta_{1},\zeta_{2}\leq 1$. In this square shape, the cell density is uniform, and cell shapes are isotropic. Next, we need to specify the active tension configuration, i.e., the tension metric $g$. In this example, we consider an anisotropic metric, $g_{\zeta\zeta}=2,\quad g_{\zeta\bar{\zeta}}=s$ where the real parameter $0<s\ll 1$ determines the strength of the anisotropy. In Cartesian notation, $\mathbf{g}(\zeta_{1},\zeta_{2})=\mathrm{diag}[1+s,1-s]$. Microscopically, this tension metric means that vertical cell-cell interfaces are under higher tension than horizontal ones, since the tension metric encodes the tension on interfaces transverse to a line element $d\mathbf{r}$ (cf. Fig. 2). The Beltrami coefficient reads $\mu_{g}=s/(2+\sqrt{(1+s)(1-s)})\approx s/4$, where the approximation is for $s\ll 1$. To remove this anisotropy, the $\mathbf{g}$-isothermal coordinates need to create a corresponding shear, $z(\zeta,\bar{\zeta})=\zeta+\mu_{g}\bar{\zeta}$. This fulfills the Beltrami Eq. (10). In Cartesian notation, $\mathbf{z}=\mathrm{diag}[1+\mu_{g},1-\mu_{g}]\cdot\bm{\zeta}$.
Figure 6: Elementary example for tension- and adjacency metrics.
Since $\det\mathbf{g}\approx 1$, tension and cell densities are (approximately) equal in this example. Further, since $\mathbf{g}$ is constant, it has no curvature, $K_{g}=0$: the $z$-configuration has constant cell density and constant pressure $p_{0}$. It is therefore mechanically balanced in the absence of external forces. In other words, $u_{\alpha}=0$ solves the linear elasticity force-balance equation Eq. (33). This elementary example illustrates how tension anisotropy controls tissue shape.
Let us now consider applying a constant boundary traction force $b_{1}=0,b_{2}=\pm b$ to the upper (++) and lower (-) edge of the rectangular tissue patch, (the left and right sides remain free). This results in a potential displacement $u_{\alpha}=f_{\alpha}+\partial_{\alpha}\theta$ from the tension reference $z$ to the physical configuration $w=z+u(z,\bar{z})$. By Eq. (36), the potential $\theta$ must be biharmonic, $\Delta^{2}\theta=0$. The traction forces $\sigma\cdot\mathbf{n}_{\pm}=[0,\pm b]$ at the top and bottom edges require $\sigma_{22}|_{\pm}=p_{0}(\partial_{1}^{2}\theta)|_{\pm}=b$. We can hence take $\theta=b\operatorname{Re}[z]^{2}/(2p_{0})$, which implies a uniform stress $\bm{\sigma}=\mathrm{diag}\,(0,b)$ in the bulk and, thus, also fulfills the free BCs on the left and right edge. This result is expected by simple physical reasoning. To ensure that the displacement $u_{\alpha}$ is incompressible, we add a conformal contribution f(z)=bzf(z)=-bz (in this case, simply a uniform scaling).
G.2 Statics: non-flat tension metric
Next, we consider an example where the tension network has curvature. We consider a circular tissue patch. The adjacency metric is the Euclidean metric in the unit disk. The $\mathbf{a}$-isothermal $\zeta$-coordinates are simply Cartesian coordinates. In $\zeta$-coordinates, the cell density is uniform. We now define the tension metric in the disk configuration:
We now need to find the transformation $z(\zeta,\bar{\zeta})$ that makes this tension metric isotropic. Because of rotation symmetry, we expect a radial dilation. Indeed, the $g$-isothermal coordinates are given by the stereographic projection [40] of the hemisphere, which is a conformal map:
$z(\zeta,\bar{\zeta})=\frac{\zeta}{1+h(\zeta,\bar{\zeta})}$ (84)
Due to curvature, area distortion in the isothermal coordinates, and thus cell density gradients in the tension reference state $z$ are unavoidable. For an incompressible tissue, these must be compensated by an isogonal displacement $w-z=u=f(z)+2\bar{\partial}_{z}\theta$ (Eq. (33)). Incompressibility means $(\det\mathbf{D}w)^{-1}\sqrt{\det g(z,\bar{z})}=1$ where $(Dw)_{\alpha\beta}=\partial_{\alpha}w_{\beta}$ is the Jacobian. Expanding $\det\mathbf{D}w=\det[\delta_{\alpha\beta}+\partial_{\alpha}u_{\beta}]$ with $w_{\alpha}=f_{\alpha}+\partial_{\alpha}\theta$ to linear order, this implies
$\Delta\theta+\partial_{\alpha}f_{\alpha}=\frac{2}{z\bar{z}-1}$ (85)
$\Delta\theta(r)+C=\frac{2}{1+r^{2}}$ (86)
with BCs $\theta(0)=0$, $\partial_{r}\theta(r=0)=0$, and $\partial_{r}\theta(r=0)=0$. The last condition follows from $\sigma\cdot\mathbf{n}=0$ (vanishing traction forces at the boundary). The strain $C=\partial_{\alpha}f_{\alpha}$ of the conformal map is chosen to fulfill the last BC. Eq. (86) is solved by $\theta(r)=\tfrac{1}{4}[-r^{2}+2\log(1+r^{2})]$ with $C=1$.
Within the disk, $\sigma_{\alpha\beta}^{\mathrm{tot}}=-p_{0}\epsilon_{ac}\epsilon_{bd}\partial_{\gamma}\partial_{\delta}\theta\neq 0$ (Eq. (35)). These internal residual stresses, shown in Fig. 7, result from curvature of the tension manifold. Importantly, such stresses will also be present for finite cell compressibility. In the limit of fully compressible cells, no force-balanced configuration exists because sustaining the pressure field implied by the tension manifold curvature requires a finite cell compressibility.
G.3 Dynamics: positive tension feedback
We now consider an example of tissue flow driven by tension dynamics. We translate the results of Ref. [30], which analyzed a model for positive tension feedback with mean-field techniques, into the notation of the present manuscript. For simplicity, we consider the case where $w\approx z$, i.e., no tension-cell density incompatibility or external forces.
In the mean-field setting, a tissue patch is described by the local distribution of tension triangles (correlations between adjacent triangles are neglected), each of which is characterized by a Beltrami coefficient $\mu_{\zeta|z}^{\triangle}$. The macroscopic order parameters are moments of this distribution. The mean $\langle\mu_{\zeta|z}^{\triangle}\rangle=\mu_{\zeta|z}$ defines the macroscopic tension anisotropy, while the second moment $\langle|\mu_{\zeta|z}^{\triangle}|^{2}\rangle=q^{2}$ measures how anisotropic individual triangles are. Note that $q>|\mu_{\zeta|z}|$: e.g., in a disordered triangulation, triangles can be anisotropic with no net tissue-scale anisotropy. (In the notation of Ref. [30], $\mathbf{N}=\mu_{\zeta|z}/q$, neglecting correlations between $\arg\mu_{\zeta|z}^{\triangle}$ and $|\mu_{\zeta|z}^{\triangle}|$). In general, a wider distribution of tension anisotropies leads to a higher T1 rate, since there are more triangles “close” to the T1 threshold.
The parameter $q$ determines the T1 rate $\gamma_{\mathrm{T1}}(q)$: the higher $q$, the higher the rate of T1s, with $\gamma_{\mathrm{T1}}\mapsto\infty$ as $q\mapsto 1$. The dynamics of the tension anisotropy $\mu_{\zeta|z}$ combined two effects: amplification due to positive feedback with a characteristic timescale $t_{\tau}$ and degradation of anisotropy by T1s:
$\partial_{t}\mu_{\zeta|z}=\frac{1}{t_{\tau}}\mu_{\zeta|z}-\alpha(q)\gamma_{\mathrm{T1}}(q)\mu_{\zeta|z}.$ (87)
The coefficient $\alpha$ can be calculated from the mean field theory. One finds that the combined coefficient $1-\alpha\gamma_{\mathrm{T1}}t_{\tau}<0$ at long times, so that the dynamics eventually erase any tension anisotropy. However, the transient remodeling of the adjacency metric, due to the Beltrami flow $\bar{\partial}_{\xi}(\partial_{t}\zeta)=\gamma_{\mathrm{T1}}\mu_{z}$, leads to a net plastic tissue deformation due to the initial tension anisotropy.

### v3

@@ Appendix G Kinematics of quasiconformal flow
Quasiconformal flow induces $\mu_{\zeta|\zeta^{\prime}}=\bar{\partial}_{\zeta}{\dot{\zeta}}dt/(1+\partial_{\zeta}{\dot{\zeta}}dt)\approx\bar{\partial}_{\zeta}{\dot{\zeta}}dt=\partial_{t}\mu_{a}dt$. To find the Beltrami coefficient of the inverse of a map, one applies the chain rule:
@@
Now composing with the map $\zeta\mapsto z$ gives
@@
where the phase factor is $e^{-i2\phi_{\zeta^{\prime}|\zeta}}=(1-\overline{\partial_{\zeta}{\dot{\zeta}}}dt)/(1-\partial_{\zeta}{\dot{\zeta}}dt)=1+i2\mathrm{Im}[\partial_{\zeta}{\dot{\zeta}}]dt$. The displacement change from $\zeta$ to $\zeta^{\prime}$ leads to an additional advective term $\mathrm{Re}[\bar{\nu}\partial_{\zeta}]\mu_{z}dt$. Taken together, one arrives at
@@
$=\partial_{t}\mu_{g}-\partial_{t}\mu_{a}+\mathrm{Re}[\dot{\zeta}\partial_{\zeta}]\mu_{\zeta|z}$
@@
The final, nonlinear term can be neglected for weak anisotropy $|\mu_{\zeta|z}|\ll 1$ and captures strain-induced rotation. In conventional Cartesian notation, the total derivative of the tension metric $\mathbf{g}$ can be written as
$\frac{d\mathbf{g}(\zeta,\bar{\zeta})}{dt}=\mathbf{m}+\left[({\dot{\zeta}}\cdot\mathbf{D})\mathbf{g}-\mathbf{D}{\dot{\zeta}}\cdot\mathbf{g}-\mathbf{g}\cdot(\mathbf{D}{\dot{\zeta}})^{T}\right]$ (90)
Figure 8 provides a geometric explanation of Eq. (88). The term $\partial_{t}\mu_{g}$ arises from the intrinsic dynamics of $z$, in addition to the coordinate transformation $\zeta\mapsto\zeta^{\prime}$.
G.2 Variation of $|\mu_{\zeta|w}|^{2}$
@@
makes clear that the Beltrami coefficient of the new map $\zeta^{\prime}\mapsto w$ is calculated by composing $\zeta^{\prime}\mapsto\zeta\mapsto w$ [red path in Eq. (91)]. We first use that $\zeta\mapsto\zeta^{\prime}$ is simply the quasiconformal flow induced by $\delta\mathbf{a}$ so $\mu_{\zeta|\zeta^{\prime}}=\bar{\partial}_{\zeta}\zeta^{\prime}/\partial_{\zeta}\zeta^{\prime}=\delta\mu_{a}$. The inverse is obtained using Eq. (86), giving